6.895 Theory of Parallel Systems Seth Gilbert

Problems 3–5

Problem 3. Let us pretend that the first problem on this problem set, #3, involves explaining how to write-up problems using the 6895.sty latex file. In this case, begin the document normally, and use the 6895.sty package. The next step, is to begin the problems. This is accomplished by:

     \begin{problems}{name}{first problem num}{last problem num}

At this point, we are ready to actually write-up problems. You the following problem environment:

     \begin{problem}
     \end{problem}

You can also do subproblems similarly:

     \begin{problempart}
     \end{problempart}

To reset the subcounter counter, either embed the subproblem environment within a problem environment, or use:

     \resetsubprobs

Problem 4. Now we pretend that problem 4 is advice on writing up problem sets. First, see the advice on writing scribe notes. As in scribe notes, use theorem/proof environments to make clear what you are trying to prove (and which part is actually the proof, and not just an overview). This is important. Otherwise, it will be quite hard for me to determine what you have proved and how you are proving it! Similarly, use all the other nice macros in 6895.sty. Also, it is recommended that every problem include a figure. This makes it much easier to understand what you are trying to prove (and might even help you understand how to prove it.) This can be drawn with xfig, or handed in separately.

Problem 5. Now we pretend that problem 5 is more instructions on problems. In order to submit a problem, e-mail the latex source and postscript to sethg@mit.edu. The files should have a unique name. That is, name the files: yourname uniquenumber.tex and yourname uniquenumber.ps. If you are in Singapore, e-mail these to the Singapore TA. Any other hardcopy materials should be left at my office NE43-371 (or in Singapore at the TAs office).